# Surrogate-based multilevel Monte Carlo methods for uncertainty quantification in the Grad-Shafranov free boundary problem

### v1

@@ 3.2 Sampling with surrogates with multilevel spatial discretization
$N_{\ell}=\left\lceil\frac{1}{\theta\epsilon^{2}\left\|\mathbb{E}(u_{L})\right\|_{Z}^{2}}\sqrt{\frac{\mathbb{V}\left(\widehat{Y}_{\ell}\right)}{W_{\ell}^{e}}}\sum_{k=0}^{L}\sqrt{\mathbb{V}\left(\widehat{Y}_{k}\right)W_{k}^{e}}\right\rceil.$ (32)
There is, however, one issue with this estimate. For any given level $\ell$, the expression above requires knowledge of the variance $\mathbb{V}\left(\widehat{Y}_{k}\right)$ for all levels $0\leq k\leq L$. In turn, computing or approximating the variance, for instance through the sample variance
$\mathbb{V}\left(\widehat{Y}_{\ell}\right)=\frac{1}{N_{\ell}-1}\left(\sum_{i=1}^{N_{\ell}}\left\|\widehat{Y}_{\ell}^{(i)}\right\|_{Z}^{2}-\frac{1}{N_{\ell}}\left\|\sum_{i=1}^{N_{\ell}}\widehat{Y}_{\ell}^{(i)}\right\|_{Z}^{2}\right)\,,$ (33)
@@
$\mathbb{V}(u-\widehat{u}_{\ell})=\mathbb{E}\left[\left\|u-\widehat{u}_{\ell}\right\|_{Z}^{2}\right]-\left\|\mathbb{E}\left(u-\widehat{u}_{\ell}\right)\right\|_{Z}^{2}\leq\mathbb{E}\left[\left\|u-\widehat{u}_{\ell}\right\|_{Z}^{2}\right]=\|u-\widehat{u}_{L}\|^{2}_{L^{2}(\boldsymbol{W},Z)}.$
As shown in (2.4), by selecting the number of sparse grid nodes according to (20) and from the bounds (11) and (13), the total error $\|u-\widehat{u}_{L}\|_{L^{2}(\boldsymbol{W},Z)}$ decays at the same rate $\alpha$ as the discretization error, which serves as an upper bound for $\mathbb{V}(u-\widehat{u}_{\ell})$. By virtue of the subtraction in the equation above, this leads to
$\mathbb{V}(u-\widehat{u}_{\ell})\simeq M_{\ell}^{-\beta_{1}},$ (34)
where $\beta_{1}\geq 2\alpha$. Let us also assume that the variance of the difference between surrogates at two successive levels satisfies a similar bound, i.e. $\mathbb{V}(\widehat{u}_{\ell+1}-\widehat{u}_{\ell})\simeq M_{\ell}^{-\beta_{1}}$. With $\mathbb{V}(\widehat{Y}_{\ell+1})=\mathbb{V}(\widehat{u}_{\ell+1}-\widehat{u}_{\ell})$, and using the mesh growth assumption (15), it then follows that
$\mathbb{V}\left(\widehat{Y}_{\ell+1}\right)\propto(M_{\ell}/M_{\ell-1})^{-\beta_{1}}\,\mathbb{V}\left(\widehat{Y}_{\ell}\right)=s^{-\beta_{1}}\mathbb{V}\left(\widehat{Y}_{\ell}\right).$ (35)
This extrapolation estimate allows us to approximate the variance of subsequent levels based on the sample variances of available levels, giving rise to the following iterative method to estimate the number of required samples at each level. Starting with an initial guess $N_{0}$ for the number of samples at the coarsest level, the sample variance $\mathbb{V}(\widehat{Y}_{0})$ is computed, and then for any subsequent level $\ell$ we extrapolate $\mathbb{V}(\widehat{Y}_{\ell+1})$ from $\mathbb{V}(\widehat{Y}_{\ell})$ using (35). The extrapolated values are then substituted into (32) and the prescribed number of samples is gathered for every discretization level. With these samples the sample variances are computed and a measure of the statistical error, as implicitly defined in the final term of (29), is obtained. If the approximation condition (30) for the statistical error is satisfied, the algorithm stops. Otherwise the number of samples is recomputed using the updated values of the variance and the process is repeated until condition (30) is met. The process is described in Algorithm 1. This algorithm is derived from Giles’ basic theorem [15] and it accounts for both discretization and statistical errors. A key difference is that the algorithm uses a surrogate model to evaluate samples, with the interpolation error reflected through the construction of a surrogate in the algorithm input.
@@
Output: Fine-grid level $L$, vector of sample sizes $\boldsymbol{N}=(N_{0}\ldots,N_{L})$ for each discretization level, vector of variance estimations $\boldsymbol{V}=(\mathbb{V}(\widehat{Y}_{0}),\ldots,\mathbb{V}(\widehat{Y}_{L}))$, expectation estimate $A$.
@@
5Initialize vector of sample number corrections $\boldsymbol{dN}\leftarrow\boldsymbol{N}_{\text{old}}$ while $\sum_{\ell}dN_{\ell}>0$ do
@@
10
11      Approximate $\mathbb{V}(\widehat{Y}_{\ell})$ for $0\leq\ell\leq L$ using (33).
@@
16
17            if  $\left\|u-u_{L}\right\|_{L^{2}(\boldsymbol{W},Z)}<\theta_{1}\epsilon\left\|\mathbb{E}(u)\right\|_{Z}$, then
18                   Compute $A(\widehat{u}_{L})$ by (28).
@@
21                  Approximate $\mathbb{V}(\widehat{Y}_{L})$ by (35).
@@
$\sum_{\ell=0}^{L}N_{\ell}W_{\ell}^{e}\geq\sum_{\ell=0}^{L}\left(\frac{1}{\theta\epsilon^{2}\left\|\mathbb{E}(u_{L})\right\|_{Z}^{2}}\sqrt{\frac{\mathbb{V}\left(\widehat{Y}_{\ell}\right)}{W_{\ell}^{e}}}\sum_{k=0}^{L}\sqrt{\mathbb{V}\left(\widehat{Y}_{k}\right)W_{k}^{e}}\right)W_{\ell}^{e}=\frac{1}{\theta\epsilon^{2}\left\|\mathbb{E}(u_{L})\right\|_{Z}^{2}}\left(\sum_{\ell=0}^{L}\sqrt{\mathbb{V}\left(\widehat{Y}_{\ell}\right)W_{\ell}^{e}}\right)^{2}.$
@@
$\sum_{\ell=0}^{L}N_{\ell}W_{\ell}^{e}\geq\frac{\mathbb{V}\left(\widehat{Y}_{L}\right)W_{L}^{e}}{\theta\epsilon^{2}\left\|\mathbb{E}(u_{L})\right\|_{Z}^{2}},$ (36)
@@
$\mathbb{V}\left(\widehat{Y}_{L}\right)=s^{-\beta_{1}L}\mathbb{V}\left(\widehat{Y}_{0}\right)=s^{-\beta_{1}L}\mathbb{V}\left(\widehat{u}_{0}\right),$ (38)
@@
$\mathbb{V}\left(\widehat{Y}_{L}\right)=s^{-\beta_{1}L}\mathbb{V}\left(\widehat{u}_{0}\right)\simeq s^{-\frac{\beta_{1}}{\alpha}\log_{s}\left(\epsilon^{-1}\right)}\simeq\epsilon^{\frac{\beta_{1}}{\alpha}}.$
@@
$\sum_{\ell=0}^{L}N_{\ell}W_{\ell}^{e}\geq c(\epsilon),\quad\mbox{where}\quad c(\epsilon)\simeq\left(\frac{1}{\theta\epsilon^{2}\left\|\mathbb{E}(u_{L})\right\|_{Z}^{2}}\right)\epsilon^{2}\epsilon^{-\frac{1}{\alpha}-\frac{\delta}{\nu}}\simeq\epsilon^{-\frac{1}{\alpha}-\frac{\delta}{\nu}}.$
@@
$\sum_{\ell=0}^{L}N_{\ell}W_{\ell}^{e}\leq\sum_{\ell=0}^{L}\left(\frac{2}{\theta\epsilon^{2}\left\|\mathbb{E}(u_{L})\right\|_{Z}^{2}}\sqrt{\frac{\mathbb{V}\left(\widehat{Y}_{\ell}\right)}{W_{\ell}^{e}}}\sum_{k=0}^{L}\sqrt{\mathbb{V}\left(\widehat{Y}_{k}\right)W_{k}^{e}}\right)W_{\ell}^{e}=\frac{2}{\theta\epsilon^{2}\left\|\mathbb{E}(u_{L})\right\|_{Z}^{2}}\left(\sum_{\ell=0}^{L}\sqrt{\mathbb{V}\left(\widehat{Y}_{k}\right)W_{k}^{e}}\,\right)^{2}.$
We now focus our attention on the term $\sum_{k=0}^{L}\sqrt{\mathbb{V}\left(\widehat{Y}_{k}\right)W_{k}^{e}}$ and note that from (15) and (20) it follows that
@@
$\sum_{k=0}^{L}\sqrt{\mathbb{V}\left(\widehat{Y}_{k}\right)W_{k}^{e}}\simeq s^{L\alpha\delta/2\nu}\sum_{k=0}^{L}s^{k(1-\beta_{1})/2}\simeq\epsilon^{-\delta/2\nu}\sum_{k=0}^{L}s^{k(1-\beta_{1})/2}.$
@@ 4 Numerical experiments
We now present numerical results for surrogate-based sampling to estimate the expected poloidal flux $\mathbb{E}(u)$. We build surrogates through sparse grid collocation and explore the construction cost. We then perform MC-FE and MLMC-FE sampling with both direct computation and surrogates, assessing efficiency by comparing the computational cost, as measured by CPU times, and also evaluating the accuracy of certain quantities such as the separatrix and various geometric descriptors derived from the approximation of $\mathbb{E}(u)$. We note at the outset that in previous work [11], we found that the interpolation error appears to satisfy a bound like (7) but not (6), so that the cost analysis of the previous sections is not directly applicable. Despite this, the results suggest that collocation and multilevel MC lead to cost savings along the lines suggested by the analysis.
@@ 4.1 Experiment description
MLMC-FE sampling requires a hierarchical set of spatial meshes. Starting from a reference mesh provided by FEEQS.M, we generate this set using the Triangle mesh generator [33] to refine and derefine the reference mesh within the region enclosed by the (spline-approximated) boundaries. This process results in a collection of non-nested and geometry-conforming uniform meshes [12]. The number of grid points $M_{\ell}$ of each mesh in the hierarchy is detailed in Table 1, where the reference mesh (level 2 in the hierarchy) contains 30,449 grid points. We generate surrogates on these spatial meshes using stochastic collocation on Clenshaw-Curtis sparse grids within a 12-dimensional parameter space. The number of sparse grid nodes $P_{q}$ at increasing levels is shown in Table 1. For all sampling conducted in this study, the tolerances for the normalized mean squared error range from $\epsilon=2\times 10^{-4}$ to $8\times 10^{-3}$. The sample mean and the normalized sample variance $V_{\ell}\equiv\mathbb{V}(\widehat{Y}_{\ell})/\left\|\mathbb{E}(u_{L})\right\|_{Z}^{2}$ are dynamically updated using Welford’s algorithm
@@
For direct non-linear solves at each parameter value, the stopping threshold for the relative residual of Newton’s method is set to $5\times 10^{-11}$. The normalization factor $\|\mathbb{E}(u_{\ell})\|_{Z}$ on the finest spatial mesh level ($\ell=5$) is approximately $8.5708\times 10^{-1}$. All experiments are conducted using MATLAB R2023a on a System 76 Thelio Major with 256GB RAM and a 64-Core @4.6 GHz AMD Threadripper 3 processor.
@@ 4.3 Sampling (online) costs
We now use these surrogates to run online simulations with MC-FE and MLMC-FE sampling. For each sampling approach, we will examine the sampling cost, calculate the statistical estimator for the expected solution (3), estimate the sample size, and compare the combined offline and online costs. Note that it follows from (5) that the average cost to evaluate the surrogate on a grid with $M_{\ell}$ points is linear in $M_{\ell}$. It can also be shown that the dependence on the number of sparse grid nodes is proportional to $M_{\ell}P_{q}^{\delta}$ [24, Section 3.3.5.4]. We have found experimentally that $\delta\approx 1$ [26].
@@
Since $\gamma\approx 1$, the solid and dotted blue lines in the plot are almost parallel. [^2] Note that the value $\gamma=1.09$ was determined from experiments. In general, for large $M$, we expect $\gamma$ to be larger for sparse direct solvers of the Jacobian systems [7, 17], such as $\gamma=1.5$ for the nested dissection method [14] and closer to $2$ for a band solver [17, Section 4.3]. The cost of the methodology used for direct solution (solid blue curve) is influenced by various factors, including the cost of solving the Jacobian system, the number of iterations required by Newton’s method, complexities from the nonlinearity of the free plasma boundary, and the need to reassemble the nonlinear system. We believe that these factors lead to the higher cost of the direct solution indicated in the figure. If $\gamma$ were larger, the advantages of the surrogate solution (dashed blue curve) would be more dramatic.
@@
MC-FE Direct solver MC-FE Surrogate MLMC-FE Direct solver MLMC-FE Surrogate MLMC-FE DS Interp MLMC-FE Surrogate Interp
$\epsilon$ Time
@@
Table 4: CPU times in seconds together with speedups for the multilevel methods, for a variety of choices of $\epsilon$. For MC-FE with direct solver and surrogate, MLMC-FE with direct solver and surrogate, and MLMC-FE with direct solver and surrogate with solution interpolating to a common fine grid of level $\ell=5$. The computational cost associated with a tolerance of $\epsilon=2\times 10^{-4}$ for Monte Carlo was prohibitive; the entry in the table for this tolerance (with an asterisk) is an estimate.
@@
Level $\ell$
$\epsilon$ 0 1 2 3 4 5
$8\times 10^{-3}$ 5
$6\times 10^{-3}$ 7
$4\times 10^{-3}$ 22
$2\times 10^{-3}$ 83
$10^{-3}$ 322
$8\times 10^{-4}$ 527
$6\times 10^{-4}$ 869
$4\times 10^{-4}$ 1980
$2\times 10^{-4}$ 8000
Level $\ell$
$\epsilon$ 0 1 2 3 4 5
$8\times 10^{-3}$ 4
$6\times 10^{-3}$ 11
$4\times 10^{-3}$ 14
$2\times 10^{-3}$ 83
$10^{-3}$ 347
$8\times 10^{-4}$ 501
$6\times 10^{-4}$ 901
$4\times 10^{-4}$ 1995
$2\times 10^{-4}$ 8326
Level $\ell$
$\epsilon$ 0 1 2 3 4 5
$8\times 10^{-3}$ 10 2 2
$6\times 10^{-3}$ 12 3 2
$4\times 10^{-3}$ 32 5 2 2
$2\times 10^{-3}$ 152 26 4 2
$10^{-3}$ 691 109 18 4 2
$8\times 10^{-4}$ 841 129 23 3 2
$6\times 10^{-4}$ 1610 231 40 8 2
$4\times 10^{-4}$ 3791 589 104 15 3
$2\times 10^{-4}$ 15859 2344 375 62 13 2
Level $\ell$
$\epsilon$ 0 1 2 3 4 5
$8\times 10^{-3}$ 10 2 2 2
$6\times 10^{-3}$ 10 3 2 2
$4\times 10^{-3}$ 37 8 2 2
$2\times 10^{-3}$ 151 22 4 2 2
$10^{-3}$ 673 134 22 5 2
$8\times 10^{-4}$ 1017 173 29 8 2
$6\times 10^{-4}$ 1867 315 51 15 4 2
$4\times 10^{-4}$ 4160 686 108 27 6 2
$2\times 10^{-4}$ 17646 2919 435 102 20 7
Table 5: Optimal sample size estimation. Left: MC-FE sampling with the direct solver. Middle left: MC-FE with the surrogate. Middle right: MLMC-FE with the direct solver. Right: MLMC-FE with the surrogate.

### v2

@@ 3.2 Sampling with surrogates with multilevel spatial discretization
$N_{\ell}=\left\lceil\frac{1}{\theta\epsilon^{2}\left\|\mathbb{E}[u_{L}]\right\|_{Z}^{2}}\sqrt{\frac{\mathbb{V}\left[\widehat{Y}_{\ell}\right]}{W_{\ell}^{e}}}\sum_{k=0}^{L}\sqrt{\mathbb{V}\left[\widehat{Y}_{k}\right]W_{k}^{e}}\right\rceil.$ (32)
There is, however, one issue with this estimate. For any given level $\ell$, the expression above requires knowledge of the variance $\mathbb{V}\left[\widehat{Y}_{k}\right]$ for all levels $0\leq k\leq L$. In turn, computing or approximating the variance, for instance through the sample variance
$\mathbb{V}\left[\widehat{Y}_{\ell}\right]=\frac{1}{N_{\ell}-1}\left(\sum_{i=1}^{N_{\ell}}\left\|\widehat{Y}_{\ell}^{(i)}\right\|_{Z}^{2}-\frac{1}{N_{\ell}}\left\|\sum_{i=1}^{N_{\ell}}\widehat{Y}_{\ell}^{(i)}\right\|_{Z}^{2}\right)\,,$ (33)
@@
$\mathbb{V}[u-\widehat{u}_{\ell}]=\mathbb{E}\left[\left\|u-\widehat{u}_{\ell}\right\|_{Z}^{2}\right]-\left\|\mathbb{E}\left[u-\widehat{u}_{\ell}\right]\right\|_{Z}^{2}\leq\mathbb{E}\left[\left\|u-\widehat{u}_{\ell}\right\|_{Z}^{2}\right]=\|u-\widehat{u}_{L}\|^{2}_{L^{2}(\boldsymbol{W},Z)}.$
As shown in (2.4), by selecting the number of sparse grid nodes according to (20) and from the bounds (11) and (13), the total error $\|u-\widehat{u}_{L}\|_{L^{2}(\boldsymbol{W},Z)}$ decays at the same rate $\alpha$ as the discretization error, which serves as an upper bound for $\mathbb{V}[u-\widehat{u}_{\ell}]$. By virtue of the subtraction in the equation above, this leads to
$\mathbb{V}[u-\widehat{u}_{\ell}]\simeq M_{\ell}^{-\beta_{1}},$ (34)
where $\beta_{1}\geq 2\alpha$. Let us also assume that the variance of the difference between surrogates at two successive levels satisfies a similar bound, i.e. $\mathbb{V}[\widehat{u}_{\ell+1}-\widehat{u}_{\ell}]\simeq M_{\ell}^{-\beta_{1}}$. With $\mathbb{V}[\widehat{Y}_{\ell+1}]=\mathbb{V}[\widehat{u}_{\ell+1}-\widehat{u}_{\ell}]$, and using the mesh growth assumption (15), it then follows that
$\mathbb{V}\left[\widehat{Y}_{\ell+1}\right]\propto(M_{\ell}/M_{\ell-1})^{-\beta_{1}}\,\mathbb{V}\left[\widehat{Y}_{\ell}\right]=s^{-\beta_{1}}\mathbb{V}\left[\widehat{Y}_{\ell}\right].$ (35)
This extrapolation estimate allows us to approximate the variance of subsequent levels based on the sample variances of available levels, giving rise to the following iterative method to estimate the number of required samples at each level. Starting with an initial guess $N_{0}$ for the number of samples at the coarsest level, the sample variance $\mathbb{V}[\widehat{Y}_{0}]$ is computed, and then for any subsequent level $\ell$ we extrapolate $\mathbb{V}[\widehat{Y}_{\ell+1}]$ from $\mathbb{V}[\widehat{Y}_{\ell}]$ using (35). The extrapolated values are then substituted into (32) and the prescribed number of samples is gathered for every discretization level. With these samples the sample variances are computed and a measure of the statistical error, as implicitly defined in the final term of (29), is obtained. If the approximation condition (30) for the statistical error is satisfied, the algorithm stops. Otherwise the number of samples is recomputed using the updated values of the variance and the process is repeated until condition (30) is met. The process is described in Algorithm 1. This algorithm is derived from Giles’ basic theorem [17] and it accounts for both discretization and statistical errors. A key difference is that the algorithm uses a surrogate model to evaluate samples, with the interpolation error reflected through the construction of a surrogate in the algorithm input.
@@
Output: Fine-grid level $L$, vector of sample sizes $\boldsymbol{N}=(N_{0}\ldots,N_{L})$ for each discretization level, vector of variance estimations $\boldsymbol{V}=(\mathbb{V}[\widehat{Y}_{0}],\ldots,\mathbb{V}[\widehat{Y}_{L}])$, expectation estimate $A$.
@@
5Initialize vector of sample number corrections $\boldsymbol{dN}\leftarrow\boldsymbol{N}_{\text{old}}$
6while $\sum_{\ell}dN_{\ell}>0$ do
@@
11
12      Approximate $\mathbb{V}[\widehat{Y}_{\ell}]$ for $0\leq\ell\leq L$ using (33).
@@
17
18            if  $\left\|u-u_{L}\right\|_{L^{2}(\boldsymbol{W},Z)}<\theta_{1}\epsilon\left\|\mathbb{E}[u]\right\|_{Z}$, then
19                   Compute $A[\widehat{u}_{L}]$ by (28).
@@
22                  Approximate $\mathbb{V}[\widehat{Y}_{L}]$ by (35).
@@
$\sum_{\ell=0}^{L}N_{\ell}W_{\ell}^{e}\geq\sum_{\ell=0}^{L}\left(\frac{1}{\theta\epsilon^{2}\left\|\mathbb{E}[u_{L}]\right\|_{Z}^{2}}\sqrt{\frac{\mathbb{V}\left[\widehat{Y}_{\ell}\right]}{W_{\ell}^{e}}}\sum_{k=0}^{L}\sqrt{\mathbb{V}\left[\widehat{Y}_{k}\right]W_{k}^{e}}\right)W_{\ell}^{e}=\frac{1}{\theta\epsilon^{2}\left\|\mathbb{E}[u_{L}]\right\|_{Z}^{2}}\left(\sum_{\ell=0}^{L}\sqrt{\mathbb{V}\left[\widehat{Y}_{\ell}\right]W_{\ell}^{e}}\right)^{2}.$
@@
$\sum_{\ell=0}^{L}N_{\ell}W_{\ell}^{e}\geq\frac{\mathbb{V}\left[\widehat{Y}_{L}\right]W_{L}^{e}}{\theta\epsilon^{2}\left\|\mathbb{E}[u_{L}]\right\|_{Z}^{2}},$ (36)
@@
$\mathbb{V}\left[\widehat{Y}_{L}\right]=s^{-\beta_{1}L}\mathbb{V}\left[\widehat{Y}_{0}\right]=s^{-\beta_{1}L}\mathbb{V}\left[\widehat{u}_{0}\right],$ (38)
@@
$\mathbb{V}\left[\widehat{Y}_{L}\right]=s^{-\beta_{1}L}\mathbb{V}\left[\widehat{u}_{0}\right]\simeq s^{-\frac{\beta_{1}}{\alpha}\log_{s}\left(\epsilon^{-1}\right)}\simeq\epsilon^{\frac{\beta_{1}}{\alpha}}.$
@@
$\sum_{\ell=0}^{L}N_{\ell}W_{\ell}^{e}\geq c(\epsilon),\quad\mbox{where}\quad c(\epsilon)\simeq\left(\frac{1}{\theta\epsilon^{2}\left\|\mathbb{E}[u_{L}]\right\|_{Z}^{2}}\right)\epsilon^{2}\epsilon^{-\frac{1}{\alpha}-\frac{\delta}{\nu}}\simeq\epsilon^{-\frac{1}{\alpha}-\frac{\delta}{\nu}}.$
@@
$\sum_{\ell=0}^{L}N_{\ell}W_{\ell}^{e}\leq\sum_{\ell=0}^{L}\left(\frac{2}{\theta\epsilon^{2}\left\|\mathbb{E}[u_{L}]\right\|_{Z}^{2}}\sqrt{\frac{\mathbb{V}\left[\widehat{Y}_{\ell}\right]}{W_{\ell}^{e}}}\sum_{k=0}^{L}\sqrt{\mathbb{V}\left[\widehat{Y}_{k}\right]W_{k}^{e}}\right)W_{\ell}^{e}=\frac{2}{\theta\epsilon^{2}\left\|\mathbb{E}[u_{L}]\right\|_{Z}^{2}}\left(\sum_{\ell=0}^{L}\sqrt{\mathbb{V}\left[\widehat{Y}_{k}\right]W_{k}^{e}}\,\right)^{2}.$
We now focus our attention on the term $\sum_{k=0}^{L}\sqrt{\mathbb{V}\left[\widehat{Y}_{k}\right]W_{k}^{e}}$ and note that from (15) and (20) it follows that
@@
$\sum_{k=0}^{L}\sqrt{\mathbb{V}\left[\widehat{Y}_{k}\right]W_{k}^{e}}\simeq s^{L\alpha\delta/2\nu}\sum_{k=0}^{L}s^{k(1-\beta_{1})/2}\simeq\epsilon^{-\delta/2\nu}\sum_{k=0}^{L}s^{k(1-\beta_{1})/2}.$
@@ 4 Numerical experiments
We now present numerical results for surrogate-based sampling to estimate the expected poloidal flux $\mathbb{E}[u]$. We build surrogates through sparse grid collocation and explore the construction cost. We then perform MC-FE and MLMC-FE sampling with both direct computation and surrogates, assessing efficiency by comparing the computational cost, as measured by CPU times, and also evaluating the accuracy of certain quantities such as the separatrix and various geometric descriptors derived from the approximation of $\mathbb{E}[u]$. We note at the outset that in previous work [13], we found that the interpolation error appears to satisfy a bound like (7) but not (6), so that the cost analysis of the previous sections is not directly applicable. Despite this, the results suggest that collocation and multilevel MC lead to cost savings along the lines suggested by the analysis.
@@ 4.1 Experiment description
MLMC-FE sampling requires a hierarchical set of spatial meshes. Starting from a reference mesh provided by FEEQS.M, we generate this set using the Triangle mesh generator [35] to refine and derefine the reference mesh within the region enclosed by the (spline-approximated) boundaries. This process results in a collection of non-nested and geometry-conforming uniform meshes [14]. The number of grid points $M_{\ell}$ of each mesh in the hierarchy is detailed in Table 1, where the reference mesh (level 2 in the hierarchy) contains 30,449 grid points. We generate surrogates on these spatial meshes using stochastic collocation on Clenshaw-Curtis sparse grids within a 12-dimensional parameter space. The number of sparse grid nodes $P_{q}$ at increasing levels is shown in Table 1. For all sampling conducted in this study, the tolerances for the normalized mean squared error range from $\epsilon=2\times 10^{-4}$ to $8\times 10^{-3}$. The sample mean and the normalized sample variance $V_{\ell}\equiv\mathbb{V}[\widehat{Y}_{\ell}]/\left\|\mathbb{E}[u_{L}]\right\|_{Z}^{2}$ are dynamically updated using Welford’s algorithm
@@
For direct non-linear solves at each parameter value, the stopping threshold for the relative residual of Newton’s method is set to $5\times 10^{-11}$. The normalization factor $\|\mathbb{E}[u_{\ell}]\|_{Z}$ on the finest spatial mesh level ($\ell=5$) is approximately $8.5708\times 10^{-1}$. All experiments are conducted using MATLAB R2023a on a System 76 Thelio Major with 256GB RAM and a 64-Core @4.6 GHz AMD Threadripper 3 processor.
@@ 4.3 Sampling (online) costs
We now use these surrogates to run online simulations with MC-FE and MLMC-FE sampling. For each sampling approach, we will examine the sampling cost, calculate the statistical estimator for the expected solution (3), estimate the sample size, and compare the combined offline and online costs. Note that it follows from (5) that the average cost to evaluate the surrogate on a grid with $M_{\ell}$ points is linear in $M_{\ell}$. It can also be shown that the dependence on the number of sparse grid nodes is proportional to $M_{\ell}P_{q}^{\delta}$ [26, Section 3.3.5.4]. We have found experimentally that $\delta\approx 1$ [28].
@@
Since $\gamma\approx 1$, the solid and dotted blue lines in the plot are almost parallel. [^2] Note that the value $\gamma=1.09$ was determined from experiments. In general, for large $M$, we expect $\gamma$ to be larger for sparse direct solvers of the Jacobian systems [9, 19], such as $\gamma=1.5$ for the nested dissection method [16] and closer to $2$ for a band solver [19, Section 4.3]. The cost of the methodology used for direct solution (solid blue curve) is influenced by various factors, including the cost of solving the Jacobian system, the number of iterations required by Newton’s method, complexities from the nonlinearity of the free plasma boundary, and the need to reassemble the nonlinear system. We believe that these factors lead to the higher cost of the direct solution indicated in the figure. If $\gamma$ were larger, the advantages of the surrogate solution (dashed blue curve) would be more dramatic.
@@
Interpolation Interpolation Heat flow Heat flow
MC-FE MC-FE MLMC-FE MLMC-FE MLMC-FE MLMC-FE MLMC-FE MLMC-FE
$\epsilon$ Direct solver Surrogate Direct solver Surrogate Direct solver Surrogate Direct solver Surrogate
Time
Time Speedup
Time Speedup
@@
4.29e+03 224.2
8.99e+01 1.1e+04
Table 4: CPU times in seconds together with speedups for the multilevel methods, for a variety of choices of $\epsilon$. For the interpolation post processing a common fine grid of level $\ell=5$ was used. The computational cost associated with a tolerance of $\epsilon=2\times 10^{-4}$ for Monte Carlo was prohibitive; the entry in the table for this tolerance (with an asterisk) is an estimate.
@@
MC-FE + Direct solver MC-FE + Surrogate MLMC-FE + Direct solver MLMC-FE + Surrogate
Level $\ell$ Level $\ell$ Level $\ell$ Level $\ell$
$\epsilon$ 0 1 2 3 4 5 0 1 2 3 4 5 0 1 2 3 4 5 0 1 2 3 4 5
$8\times 10^{-3}$ 5 4 10 2 2 10 2 2 2
$6\times 10^{-3}$ 7 11 12 3 2 10 3 2 2
$4\times 10^{-3}$ 22 14 32 5 2 2 37 8 2 2
$2\times 10^{-3}$ 83 83 152 26 4 2 151 22 4 2 2
$10^{-3}$ 322 347 691 109 18 4 2 673 134 22 5 2
$8\times 10^{-4}$ 527 501 841 129 23 3 2 1017 173 29 8 2
$6\times 10^{-4}$ 869 901 1610 231 40 8 2 1867 315 51 15 4 2
$4\times 10^{-4}$ 1980 1995 3791 589 104 15 3 4160 686 108 27 6 2
$2\times 10^{-4}$ 8000 8326 15859 2344 375 62 13 2 17646 2919 435 102 20 7
Table 5: Optimal sample size estimation. In the case of Monte Carlo estimation on the finest grid, the computational cost associated with a tolerance of $\epsilon=2\times 10^{-4}$ was prohibitive; the entry in the table for this tolerance (with an asterisk) is an estimate.